find the area of a sector if the central angle measures 5pi/6 radians and the radius of the circle is 8 centimeters. round to the nearest tenth.
Answer by RAY100(1637)

Area = pi*r^2 =pi*8^2 =64*pi
.
Area of Sector = Area * radians of sector/ 2*pi radians
.
a= ( 64 * pi) * (5/6 pi) / 2 (pi) = 26.67*pi = 83.7758= 83.8 cm ^2
.
.
check ,,,5/6 pi * 180/pi = 150 degrees, a= 64 pi * 150/360 = 83.7758 ,,,,ok
.